i have an equation 2y=5x+z where x=10 then determine y
please give me solution

Answer by Jk22(389)

y=(5x+z)/2 | divide by 2

y=(50+z)/2=25+z/2 | substitute x by 10, distribute

or
2y=50+z | substitute

y=25+z/2 | divide